OpenStudy (anonymous):

Find the slope of the line. 5x+4y=-1 Write your answer in simplest form.

OpenStudy (anonymous):

5/4x i think or negative

OpenStudy (anonymous):

\(\Huge\sf\color{#00ffff}{i got~this}\)

OpenStudy (anonymous):

Finding slope from an equation means you have to move it around until you have it in y=mx+b format. So: 5x+4y=-1 4y=-5x-1 y=-5/4x-1/4 So the slope is -5/4, hope this helped :)